FIND THE EQUATION OF THE PARABOLA WITH FOCUS (2,5) AND DIRECTRIX Y = 1
Answer by lwsshak3(11628)

FIND THE EQUATION OF THE PARABOLA WITH FOCUS (2,5) AND DIRECTRIX Y = 1
**
Standard form of equation for a parabola that opens upwards: (x-h)^2=4p(y-k), (h,k)=(x,y) coordinates of the vertex.
For given parabola:
axis of symmetry: x=2
x-coordinate of vertex=2
y-coordinate of vertex=3 (halfway between focus(5) and directrix(1) on the axis of symmetry
vertex: (2,3)
P=2 (distance from focus to vertex)
4p=8
Equation of given parabola:
(x-2)^2=8(y-3)